OpenStudy (anonymous):

If (a, -8) is a solution to the equation -a = 4b -7, what is a?

OpenStudy (anonymous):

-a = 4(-8) -7 -a = -32 -7 -a = - 39 a = 39

OpenStudy (gorv):

-a=4* (-8)-7 -a=-39 a=39